OpenStudy (anonymous):

Given the equation y − 3 =1/6 (x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. x − 2y = −12 y =1/2 x y =1/2 x + 6 y =1/2 x + 9

OpenStudy (anonymous):

so the answer wold be 54

OpenStudy (helder_edwin):

\[ \large y-3=\frac{1}{6}(x+6) \] \[ \large y-3=\frac{1}{6}x+1 \] \[ \large y=\frac{1}{6}x+4 \]

OpenStudy (anonymous):

the oones that are under the questn one of those haeve to be the answer because those are the chocies that i have to choose from

OpenStudy (helder_edwin):

there must be a typo in your question or the choices u wrote.

OpenStudy (anonymous):

oh yea i did it suppose to be y-3=1/2(x+6) sorry for that

OpenStudy (helder_edwin):

do what i did with 1/2 instead of 1/6

OpenStudy (anonymous):

yea

OpenStudy (helder_edwin):

what did u get?

OpenStudy (anonymous):

y=1/2 x+9

OpenStudy (helder_edwin):

no. look at this \[ \large y-3=\frac{1}{2}(x+6) \] \[ \large y-3=\frac{1}{2}x+3 \]

OpenStudy (anonymous):

x+6 insteead of x+9

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

thank you you were so much help

OpenStudy (helder_edwin):

u r welcome